[Hallc_running] [New Logentry] RC Update

smithg at jlab.org smithg at jlab.org
Sun Feb 23 14:55:02 EST 2020

Logentry Text:
No RC meeting today. We'll have one again tomorrow.

Run plan: continue running longitudinal at the -3.5/-3.4 GeV kinematics we returned to on Saturday. NMR every 5 hours.

Yesterday was tough. Around midday we changed to transverse as planned. But a bit over an hour into transverse running, an ioc somewhere died and it was 7 1/2 h before the associated problems in the north linac were beaten back into the bushes allowing beam delivery to resume. During this time Xiaochao determined we had enough transverse to get by, changed back to longitudinal, and setup the -3.5/-3.4 GeV kinematics during the long beam-off period. But when the next NMR was taken, the polarization had fallen from 54.9% to 43.2%. It's back up to 55.7% now. 

Overnight was reasonable, day shift too so far.

Hall access Tuesday morning, Moller cooldown tentatively Wednesday, Moller measurement tentatively Thursday.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3790388
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200223/2f8b5f67/attachment.html>

More information about the Hallc_running mailing list